1) Assume the following tax brackets are legitimate and that a client has $60,000 of economic income of which $35,000 is taxable.

Tax Bracket

Rate

0 - $10,000

10%

$10,001 - $20,000

20%

$20,001 - $50,000

30%

$50,001 and up

40%

What is the client’s tax liability?

What is the client’s ATR?

What is the client’s ETR?

What is the MTR if the client earns an additional $1 of taxable income?

2) Using the facts from example (1), assume the client earns an additional $25,000 in income, $20,000 of which is taxable.

What is the MTR?

3) Using the original facts (example 1), now assume that instead of an increase in income, your client has an unexpected deduction that reduces her taxable income by $10,000.

What is the client’s new MTR?
Answer #1

As per HOMEWORKLIB POLICY and guideline, the first-part together with all subparts is answered below:

1.

Subpart 1:

Tax liability is to be calculated on taxable income, $35,000.

Tax liability = 10,000 × 10% + (20,000 – 10,000 =) 10,000 × 20% + (35,000 – 10,000 – 10,000 =) 15,000 × 30%

                        = 1,000 + 2,000 + 4,500

                        = 7,500 (Answer)

Subpart 2:

Average tax rate (ATR) should be on taxable income.

ATR = (Tax liability / Taxable income) × 100

            = (7,500 / 35,000) × 100

            = 21.43% (Answer)

Subpart 3:

Effective tax rate (ETR) should be on economic income.

ETR = (Tax liability / Taxable income) × 100

            = (7,500 / 60,000) × 100

            = 12.50% (Answer)

Subpart 4:

Marginal tax rate (MTR) should be on the last $ income.

MTR is 30% here, since this is the latest tax rate on an additional $1 income.

MTR = 30% (Answer)